# Feedback-Free Resource Scheduling: Towards Flexible Multi-BS Cooperation in FD-RAN

### v1

@@ VI-B1 Evaluation of Proposed Greedy Algorithm
VI-B2 Evaluation of Proposed TSRA Algorithm
Under light-load network conditions, the goal is to minimize subcarrier energy cost while satisfying all user demands. Thus, the action and reward of RL need to be redefined, as the problem differs from that in heavy-load conditions. Action is defined as allocating one subcarrier from a selected combination of cooperating BSs to a user. As a result, the size of the action space is given by $|\mathcal{A}|=(2^{M}-1)\times N$. The reward should be designed to guide the agent in meeting all user demands while minimizing subcarrier energy cost. If the highest SE TM is selected, user demands can be fully met but at the cost of higher energy cost. Conversely, if the most energy-efficient TM is used to serve users, it may result in unmet user demands after all subcarriers have been allocated. Therefore, the problem involves a tradeoff between SE and EE. To this end, we introduce two types of RL to solve the problem, namely SE-prioritized and EE-prioritized RL. The reward is defined as the increase in the combined sum of SE and EE resulting from the selected action $\mathfrak{a}$. It can be expressed as $\mathfrak{r}=\beta_{SE}SE_{\mathfrak{a}}+\beta_{EE}EE_{\mathfrak{a}}$, where $\beta_{SE}$ and $\beta_{EE}$ represent the importance of SE and EE, respectively. For SE-prioritized RL, the weights are set as $\beta_{SE}=5$ and $\beta_{EE}=1$. For EE-prioritized RL, the weights are set as $\beta_{SE}=1$ and $\beta_{EE}=5$. The episode ends when either all user demands are satisfied or all subcarriers are allocated.
@@ VI-B2 Evaluation of Proposed TSRA Algorithm
Given that the second stage of TSRA may require multiple iterations to achieve optimal solutions, we evaluate its and RL’s running time over 100 test environments for networks consisting of BS1,3 and BS1,2,3 in Fig. 6. Note that the running times of SE-prioritized and EE-prioritized RL are similar, thus only EE-prioritized RL is considered in this comparison. The running time per episode for RL is relatively stable, whereas the TSRA algorithm may require additional computational time to iteratively reach optimal solutions, especially in more complex scenarios of three BSs. To ensure a fair comparison of the two algorithms’ performance, we impose a time limit on the TSRA algorithm. Using RL’s average running time as the baseline, the TSRA algorithm outputs the best solution found within the same time, denoted as "TSRA with time limits".
@@
Noticing that RL does not always satisfy all user demands, we compare the TSRA algorithm and RL in terms of the percentage of successful problem-solving cases and plot the kernel density estimation (KDE) of energy cost $e$ for all algorithms in their common successful cases. KDE is an estimation of the probability density function (PDF) of a variable using a Gaussian kernel. Fig. 7 depicts the scenario involving BS1,3. Although the TSRA algorithm is subject to a running time limit, it still quickly obtains a feasible solution, achieving a 100% success rate in solving the problem. As expected, SE-prioritized RL outperforms EE-prioritized RL in terms of problem-solving success rate, while EE-prioritized RL achieves lower energy cost. The TSRA algorithm with time constraints effectively approximates the optimal solution, demonstrating the effectiveness of the proposed approach.
@@
Compared to the two-BS scenario, three-BS cooperation increases the available TMs per user from 3 to 7, expanding the action space from 30 to 70 and making it more challenging for RL to explore and learn effectively. As illustrated in Fig. 8, although EE-prioritized RL improves the success rate of problem-solving, it tends to favor JT to meet user demands, resulting in higher energy cost. For TSRA with time constraints, although the gap from the optimal solution slightly widens in the three-BS scenario, it still achieves competitive suboptimal solutions. The subcarrier power limits demonstrate that the TSRA algorithm effectively minimizes the energy cost.
VI-B3 Comparison with Feedback-Based Resource Scheduling
@@
Then, to compare feedback-free resource scheduling with the feedback-based approach, we employ the optimal scheduling algorithm to allocate subcarriers and generate an additional 1,000 environments with random initialization. The greedy and time-constrained TSRA algorithms demonstrate superior overall performance compared to RL. Thus, they are chosen as the resource scheduling algorithms for heavy-load and light-load network conditions in subsequent comparisons. Fig. 9 presents the proportion of heavy-load and light-load network conditions for various transmission schemes across 1,000 test environments, indirectly reflecting the diversity of the dataset. In independent scheduling, where cumulative data in the buffer is not considered, the classification of the network condition as heavy-load or light-load depends on the data rate provided by the specific transmission scheme. Consequently, the presence or absence of feedback delay has little impact on network condition judgment for feedback-based schemes, as JT and the use of optimal precoders generally yield higher data rates.
(a) Without feedback delay.
(b) With feedback delay.
@@
Although the weighted sum of user demand satisfaction rate $V$ is a metric introduced under heavy-load network conditions, it should equal 1 in light-load conditions where all user demands are fully satisfied. Therefore, CDF of $V$ can be used to compare the overall scheduling performance, illustrated in Fig. 10. Optimal precoder offers better directional beams than codebooks, achieving superior performance under ideal conditions without feedback delays, as shown in Fig. 10(a). The feedback-free approach of FD-RAN leverages historical channel data to determine representative transmission parameters for each location, which results in reduced data rates compared to the ideal feedback-based approach. However, these representative transmission parameters make FD-RAN more robust in scenarios with feedback delays, allowing it to achieve the best performance among all schemes, as evidenced by Fig. 10(b). On the contrary, optimal beams are so precise that they can experience greater data rate degradation than PMI in the codebook when feedback delays are present.
@@
Figure 11: Comparison between feedback-free and feedback-based resource scheduling with feedback delay across 119 test cases of light-load conditions.
Statistical results show that 119 of 1,000 test environments fall under light-load conditions for all transmission schemes. Thus, their performance is evaluated in the presence of feedback delay, as shown in Fig. 11. It can be observed that through flexible cooperation, FD-RAN can select the most suitable single BS to serve users, achieving lower subcarrier energy cost compared to the fixed transmission of BS1 with implicit feedback in most cases. Even with delays, optimal precoder feedback is expected to achieve high data rates, making single-BS transmission with explicit feedback seemingly the most energy-efficient option. However, its highly precise beams can lead to greater fluctuations in data rates, resulting in unmet user demands, as evidenced by the mean and variations of its $V$ in Fig. 11(b). In contrast, FD-RAN achieves comparable user demand satisfaction rates to CoMP of implicit feedback while maintaining lower subcarrier energy cost.
VI-C Comparison of Consecutive Scheduling Decisions
For consecutive scheduling decisions, we generate 100 test environments, each comprising 100 consecutive time slots. Untransmitted data from the previous time slots are queued in the RLC buffer for subsequent transmission. This setup enables a performance comparison under more realistic conditions, incorporating various delays configured as $d_{1}=2$ and $d_{2}=d_{3}=3$. The location/CSI reporting period and scheduling period are set to $T_{rp}=3$ and $T_{sc}=10$, respectively. Note that both delays and periods are measured in time slots. The first time slot for scheduling decisions is set to 4, meaning no data are transmitted during the initial $104+d_{2}+d_{3}=$ 10 time slots. Thus, each test environment consists of 9 scheduling decisions. In what follows, we simulate two different values of $\rho$ to represent low and high mean data arrival rates for users: $\rho=2000$ and $\rho=4000$.
@@ VI-C Comparison of Consecutive Scheduling Decisions
(b) Subcarrier energy cost and $V$.
Figure 12: Visualization of trends of variables in FD-RAN during data transmission in the third test environment from the network’s perspective.
Figure 13: Visualization of trends of variables in FD-RAN during data transmission in the third test environment from user 9’s perspective.
@@
VI-C1 Low Mean Data Arrival Rate
First, we compare feedback-free and feedback-based resource allocation under a low mean data arrival rate at different users’ buffers. Fig. 12 and Fig. 13 illustrate the trends of variables during the data transmission in one test environment when the transmission scheme is FD-RAN, from the perspectives of the entire network and user 9, respectively. The former figure presents the network’s transitions between heavy-load and light-load conditions over consecutive time slots, as well as the fluctuating decline in the total data amount in user buffers. The latter highlights FD-RAN’s flexibility in selecting serving BSs to meet user demands. In consecutive scheduling decisions, $V$ is defined as the ratio of the total transmitted data during the scheduling period to the cumulative data preceding it. The network may transition to the heavy-load condition as data accumulates in the buffer when the selected transmission scheme provides insufficient data rates. Fig. 14 visualizes the proportion of heavy-load and light-load network conditions for various transmission schemes across 900 scheduling decisions in all test environments. CoMP with explicit feedback and FD-RAN with flexible cooperation can provide higher data rates, helping to maintain the network in light-load conditions.
@@ VI-C1 Low Mean Data Arrival Rate
Figure 16: Comparison of $e$ between FD-RAN and CoMP of explicit feedback with a low mean data arrival rate under consecutive scheduling decisions.
We also plot the average weighted sum of user demand satisfaction rates across 9 scheduling decisions in all test environments for various transmission schemes, as shown in Fig. 15(a). FD-RAN achieves the highest average user demand satisfaction rate, reaching 100% in most cases. This is because the feedback-free transmission provides stable data rates with representative transmission parameters, resulting in fewer occurrences of higher CQI, as shown in Fig. 15(b). Higher CQI leads to zero throughput, causing data to accumulate in the buffer. Since FD-RAN and CoMP of explicit feedback are more effective in maintaining the network in light-load conditions, their subcarrier energy costs are compared in Fig. 16. Due to the flexibility in selecting the most suitable cooperation set, FD-RAN demonstrates lower energy costs than the JT of two fixed BSs with optimal precoder feedback.
VI-C2 High Mean Data Arrival Rate
@@
When $\rho=4000$, total user demands become relatively high, making it impossible to fully satisfy them within each scheduling period. As a result, the network remains in heavy-load conditions. Nevertheless, FD-RAN still delivers the highest average user demand satisfaction rate by leveraging cooperation among all potential BSs while maintaining relatively few instances of higher CQI due to appropriate fixed transmission parameters, as shown in Fig. 17. Regardless of whether the mean data arrival rate is low or high, FD-RAN demonstrates consistently reliable performance through flexible cooperation, efficiently utilizing resources to meet diverse user demands.
@@ VII Conclusion
In this paper, we have studied how to realize feedback-free DL resource scheduling to enable multi-BS flexible cooperation in FD-RAN. A problem of joint cooperation set selection and subcarrier allocation is formulated, aiming primarily to satisfy personalized user demands. Then, the problem is decomposed based on network load conditions. In heavy-load conditions, priorities of user services and fairness are further considered, where a greedy algorithm is proposed and proves optimal. In light-load conditions, the problem is modeled as reducing subcarrier energy cost while still meeting user demands, where a two-stage resource allocation algorithm is proposed to quickly obtain a feasible solution and iteratively refine it to optimality. Simulation results based on ray-tracing channel data demonstrate the effectiveness of the proposed algorithms in achieving efficient resource utilization and low running time, and highlight the superiority of the feedback-free resource scheduling approach over CoMP in both performance and feasibility. Our future work will explore more diverse and advanced transmission methods such as NOMA and multi-user MIMO to further enhance the resource utilization flexibility of FD-RAN.

### v3

@@ VI-C1 Evaluation of Proposed Greedy Algorithm
VI-C2 Evaluation of Proposed TSRA Algorithm
@@
Given that the second stage of TSRA may require multiple iterations to achieve optimal solutions, we evaluate its and RL’s running time for networks consisting of BS1,3 and BS1,2,3 in Fig. 4. Since SE-prioritized and EE-prioritized RL exhibit comparable running times, only the EE-prioritized RL is considered in this comparison. The running time per episode for RL is relatively stable, whereas the TSRA algorithm may require additional computational time to iteratively reach optimal solutions, especially in more complex scenarios of three BSs. To ensure a fair comparison, we impose a time limit on the TSRA algorithm. Using RL’s average running time as the baseline, the TSRA algorithm outputs the best solution found within the same runtime, denoted as “TSRA with time limits”.
@@
Since RL cannot guarantee to meet the constraint of satisfying all user demands, we compare the TSRA algorithm and RL in terms of the percentage of successful problem-solving cases and plot the kernel density estimation (KDE) of energy cost $e$ for all algorithms in their common successful cases. KDE is an estimation of the probability density function (PDF) of a variable using a Gaussian kernel. Fig. 5 depicts the scenario involving BS1,3. Although TSRA is subject to a running time limit, it still quickly obtains a feasible solution through the feasibility pump algorithm, successfully solving the problem. As expected, SE-prioritized RL outperforms EE-prioritized RL in terms of problem-solving success rate, while EE-prioritized RL achieves lower energy cost. The TSRA algorithm with time constraints effectively approximates the optimal solution and outperforms both RL methods, demonstrating its effectiveness.
@@
Compared to the two-BS scenario, three-BS cooperation expands the action space, making it more challenging for RL to explore and learn effectively. As illustrated in Fig. 6, although EE-prioritized RL improves the success rate of problem-solving, it tends to favor JT to meet user demands, resulting in higher energy cost. For TSRA under time constraints, it still yields competitive near-optimal solutions. Meanwhile, the subcarrier energy consumed by TSRA is significantly lower than the system’s maximum energy consumption, demonstrating the necessity and value of energy-efficient scheduling.
VI-D Evaluation of Feedback-Free Resource Scheduling
We further compare resource allocation in feedback-free and feedback-based systems, so as to validate the advantage of flexible cooperation in FD-RAN. The greedy and time-constrained TSRA algorithms demonstrate superior overall performance compared to RL, thus, they are chosen as the resource scheduling algorithms for heavy-load and light-load network conditions in subsequent comparisons. To implement feedback-free resource scheduling in FD-RAN, corresponding transmission parameters at each location are pre-calculated using historical channel data via variational autoencoder (VAE) [14]. To implement feedback-based resource scheduling, both implicit and explicit CSI feedback schemes are considered. The former reports a precoding matrix indicator (PMI) from a predefined codebook, while the latter reports the optimal precoder. Accordingly, these are referred to as “feedback PMI” and “feedback optimal precoder”, respectively. Considering the increasing overheads of pilots and feedback, as well as the computational complexity of joint transmission parameters for multi-BS cooperation under practical environment, feedback-based JT is restricted to two BSs, referred to as “CoMP”. Without loss of generality, single BS transmission and CoMP in this paper utilize BS1 and BS1,3, respectively.
Moreover, we perform evaluation under independent and consecutive scheduling decisions. To simulate time-varying channels, distinct channel samples are generated for each time slot, yielding a channel coherence time of 1 ms. While the path loss is assumed to remain constant at a specific location, different random seeds are set to reflect small-scale environmental variations, incorporating ray-tracing results into the CDL model in the static O1 scenario. In independent scheduling, each test case contains two time slots with two independent channel samples, assuming no delay in executing algorithms and activating decisions (i.e., $d_{2}=d_{3}=0$). For feedback-based schemes, CSI feedback is obtained from either the first or second channel sample to simulate a 1 ms (1 TTI) feedback delay or no delay, referred to as “with delay” and “without delay”, respectively. The scheduling decision is made for the second time slot, assuming no data accumulation. In consecutive scheduling, each test case comprises 100 consecutive time slots spanning 100 ms, with 100 distinct channel samples. Untransmitted data from the previous time slots are queued in the RLC buffer for subsequent transmission. This setup enables a performance comparison under more realistic conditions, incorporating various delays configured as $d_{1}=2$ and $d_{2}=d_{3}=3$. The location/CSI reporting period and scheduling period are set to $T_{rp}=3$ and $T_{sc}=10$, respectively. Note that both delays and periods are represented in units of time slots. User locations, data arrivals, and priority weights are randomly initialized in each test case for both independent and consecutive scheduling decisions.
@@ VI-D Evaluation of Feedback-Free Resource Scheduling
VI-D1 Comparison Under Independent Scheduling Decisions
We first compare feedback-free and feedback-based resource scheduling under independent scheduling decisions. Fig. 7 presents the proportion of heavy-load and light-load network conditions for various transmission schemes, reflecting the diversity of the test cases. In independent scheduling, where cumulative data in the buffer is not considered, the classification of the network condition as heavy-load or light-load depends on the data rate provided by the specific transmission scheme. The proportion of light-load network conditions under FD-RAN is relatively high, approaching that of CoMP with optimal precoder feedback, highlighting its strong scheduling capability in effectively meeting user demands.
(a) Without delay.
(b) With delay.
@@ VI-D1 Comparison Under Independent Scheduling Decisions
Note that $V$ should equal 1 in light-load conditions where all user demands are fully satisfied. Therefore, the CDF of $V$ can be used to compare the overall scheduling performance, illustrated in Fig. 8. In independent scheduling, $V$ is defined as the ratio of correctly transmitted data to the data arriving in the same time slot. Under ideal conditions without feedback delay, optimal precoders yield more accurate directional beams than codebooks, thereby achieving superior performance, as illustrated in Fig. 8(a). The feedback-free transmission adopted by FD-RAN exploits historical channel data to determine fixed transmission parameters for each location. While this may lead to lower data rates than feedback-based schemes without delays, it enhances robustness in the presence of delays. As shown in Fig. 8(b), such robustness allows FD-RAN to outperform other approaches when feedback delays are present. In contrast, employing optimal precoders suffers greater performance degradation than PMI when outdated CSI is used.
@@
Figure 9: Comparison between feedback-free and feedback-based resource scheduling with feedback delay under light-load network conditions.
When delving into the test data, we find some test cases fall under light-load conditions for all transmission schemes. Thus, their EE performance is evaluated in the presence of feedback delay, as shown in Fig. 9. It is observed that FD-RAN, enabled by flexible cooperation, can dynamically select the most appropriate BS for each user, thereby achieving lower subcarrier energy cost than single-BS transmission with PMI feedback in most cases. Although explicit feedback with optimal precoders theoretically enables higher data rates, the high beam precision renders it more sensitive to feedback delays, resulting in greater rate fluctuations and unmet user demands. This is reflected in both the mean and variations of $V$ in Fig. 9(b). In contrast, FD-RAN achieves comparable user demand satisfaction rates to PMI-based CoMP with lower subcarrier energy cost, due to its insensitivity to delay and flexibility in selecting the most energy-efficient single-BS transmission.
VI-D2 Comparison Under Consecutive Scheduling Decisions
For consecutive scheduling decisions, we consider both low and high data arrival rates at user buffers to simulate different network load conditions.
@@ VI-D2 Comparison Under Consecutive Scheduling Decisions
(b) Subcarrier energy cost.
Figure 10: Visualization of trends of variables in FD-RAN during data transmission from the perspective of all users.
Figure 11: Visualization of trends of variables in FD-RAN during data transmission from the perspective of one user.
@@
Figure 14: Comparison of $e$ between FD-RAN and CoMP with a low mean data arrival rate under consecutive scheduling decisions.
@@
First, we compare feedback-free and feedback-based resource allocation under a low mean data arrival rate in different users’ buffers. Fig. 10 and Fig. 11 illustrate the trends of variables during the data transmission when the transmission scheme is FD-RAN, from the perspectives of all users and one user in the network, respectively. The former figure presents the network’s transitions between heavy-load and light-load conditions over consecutive time slots, as well as the fluctuation of the total data amount in user buffers. The latter highlights FD-RAN’s flexibility in selecting serving BSs to meet user demands. In consecutive scheduling decisions, the network may transition to the heavy-load condition as data accumulates in the buffer when the selected transmission scheme provides insufficient data rates. Fig. 12 visualizes the proportion of heavy-load and light-load network conditions for various transmission schemes under consecutive scheduling decisions. CoMP with optimal precoder feedback and FD-RAN with flexible cooperation can provide higher data rates, helping to maintain the network in light-load conditions.
We also plot the average $V$ for various transmission schemes, as shown in Fig. 13(a). In consecutive scheduling, $V$ is defined as the ratio of the total transmitted data during the scheduling period to the cumulative data preceding it. FD-RAN achieves the highest average $V$, reaching 1 in most cases. This is because the feedback-free transmission provides stable data rates with representative transmission parameters, resulting in fewer occurrences of bad CQI, as shown in Fig. 13(b). Bad CQI refers to the selection of a higher CQI that results in zero throughput and causes data to accumulate in the buffer. Since FD-RAN and CoMP with optimal precoder feedback are more effective in maintaining the network in light-load conditions, their subcarrier energy costs are compared in Fig. 14. Due to the flexibility in selecting the most suitable cooperation set, FD-RAN demonstrates lower energy costs.
When the mean data arrival rate at user buffers is high, total user demands exceed the network’s capacity, making it impossible to fully satisfy them within each scheduling period. As a result, the network remains in heavy-load conditions. Nevertheless, FD-RAN still delivers the highest average $V$ by leveraging cooperation among all potential BSs while maintaining relatively few instances of bad CQI due to appropriate fixed transmission parameters, as shown in Fig. 15. Regardless of whether the mean data arrival rate is low or high, FD-RAN demonstrates consistently reliable performance through flexible cooperation, efficiently utilizing resources to meet diverse user demands.
@@ VII Conclusion
In this paper, we have studied how to realize feedback-free DL resource scheduling to enable multi-BS flexible cooperation in FD-RAN. A problem of joint cooperation set selection and subcarrier allocation is formulated, aiming primarily to satisfy personalized user demands. Then, the problem is decomposed based on network load conditions. In heavy-load conditions, priorities of user services and fairness are further considered, where a greedy algorithm is proposed and proves optimal. In light-load conditions, the problem is modeled as reducing subcarrier energy cost while still meeting user demands, where a TSRA algorithm is proposed to obtain a near-optimal solution within a short time. Simulation results based on ray-tracing channel data demonstrate that the proposed algorithms achieve higher resource utilization efficiency than RL algorithms under the same runtime. Moreover, the feedback-free resource scheduling approach outperforms the two-BS CoMP scheme. Our future work will explore more diverse and advanced transmission methods such as NOMA and multi-user MIMO to further enhance the resource utilization flexibility of FD-RAN.